solve 5-x=3(2-x)+2x
Answer by funmath(2933)

solve 5-x=3(2-x)+2x
5-x=3(2)+3(-x)+2x
5-x=6-3x+2x
5-x=6-x
5-x+x=6-x+x
5=6
Your variable was eliminated and 5 and 6 are not equal so this is inconsistant and there is no solution.
Happy Calculating!!!